# A Dataset and Benchmark for Robotic Cloth Unfolding Grasp Selection: The ICRA 2024 Cloth Competition

### v1

@@ Abstract
Robotic cloth manipulation suffers from a lack of standardized benchmarks and shared datasets for evaluating and comparing different approaches. To address this, we created a benchmark and organized the ICRA 2024 Cloth Competition, a unique head-to-head evaluation focused on grasp pose selection for in-air robotic cloth unfolding. Eleven diverse teams participated in the competition, utilizing our publicly released dataset of real-world robotic cloth unfolding attempts and a variety of methods to design their unfolding approaches. Afterwards, we also expanded our dataset with 176 competition evaluation trials, resulting in a dataset of 679 unfolding demonstrations across 34 garments. Analysis of the competition results revealed insights about the trade-off between grasp success and coverage, the surprisingly strong achievements of hand-engineered methods and a significant discrepancy between competition performance and prior work, underscoring the importance of independent, out-of-the-lab evaluation in robotic cloth manipulation. The associated dataset is a valuable resource for developing and evaluating grasp selection methods, particularly for learning-based approaches. We hope that our benchmark, dataset and competition results can serve as a foundation for future benchmarks and drive further progress in data-driven robotic cloth manipulation.
@@ 2.1.1 Overview of Cloth Unfolding Strategies
Cloth unfolding is a critical prerequisite for various downstream cloth manipulation tasks such as folding, hanging, ironing, and assisted dressing. Researchers have explored various unfolding strategies, with computer vision often playing an important role. For a general introduction to using computer vision for cloth manipulation, we refer the reader to the surveys by Jiménez (2017) and Jiménez and Torras (2020). The strategy evaluated in our benchmark and competition aligns with what Jimenez identifies as Grasp Point Localisation, one of three key visual tasks he proposes as fundamental to cloth manipulation, alongside classification and state estimation. Grasp point localisation is deemed fundamental because, as Jiménez (2017) aptly states: “Any manipulative action requires instructing the robot where to grasp.” Combining well-placed grasps with open-loop motions can be sufficient for completing many manipulation tasks. For instance, unfolding might be achieved with a predetermined stretch motion after determining two appropriate grasp poses on the cloth.
@@
  • In-air grasping: Grasping a new point in the air while holding the cloth with one gripper and then stretching the cloth to unfold (Maitin-Shepard et al., 2010; Doumanoglou et al., 2014a). Starting from a crumpled garment, a common heuristic to bootstrap this strategy is first grasping its highest (or a random) point, lifting it, and then grasping its lowest point. This is the strategy used in our benchmark, and we also use the lowest point heuristic as the initialisation procedure.
  • Lifting and laying down: Simultaneous grasping of two points on a cloth lying on a surface, followed by lifting and laying it back down (Martínez and Ruiz-del Solar, 2017; Thananjeyan et al., 2022). FlingBot exemplifies this approach, incorporating a flinging motion before laying down the cloth (Ha and Song, 2022).
@@
While not mutually exclusive (humans often seamlessly switch between strategies), these categories offer a framework for understanding the diverse approaches in the literature. A key commonality among the first three categories is the focus on achieving two “good” grasps on the cloth as a primary means of unfolding. These strategies operate under the generally valid assumption that grasping a cloth item at two suitable points can largely accomplish unfolding. Furthermore, due to the nature of the unfolding task, nearly all the aforementioned strategies can be applied iteratively, with the option to reset by simply releasing the cloth and starting over. This iterative nature necessitates that evaluation metrics like success rate or final coverage are always contextualized in total amount of time taken, as even very inefficient strategies can eventually succeed through sufficient attempts.
@@ 2.2 Datasets for Robotic Cloth Manipulation
Datasets like Dex-Net 2.0 (Mahler et al., 2017) and GraspNet-1Billion (Fang et al., 2020a) were instrumental in advancing learning-based grasping, primarily for pick-and-place tasks with rigid objects. These works relied on analytical measures to determine grasp stability, focusing on feasibility and force closure. However, grasp selection for cloth unfolding presents a more complex challenge. The deformable nature of cloth makes it difficult to analytically determine grasp success, as the interaction between the gripper and the fabric can lead to unpredictable deformations. Additionally, the relative grasp location on the cloth significantly influences the unfolding outcome, making it crucial to consider not only grasp stability but also the subsequent manipulation trajectory. This distinction highlights the need for specialized cloth manipulation datasets.
@@ 3.2 Grasp Execution
To maximize both grasp success and workspace coverage while avoiding premature collisions, a two-part motion strategy is used for executing the provided grasps. This strategy consists of first moving the arm to a collision-free pregrasp pose, positioned a few centimeters behind the final grasp pose (i.e. the orientation is the same but there is a small translation along the negative z-axis of the gripper, compared to the actual grasp frame). Then, from this pregrasp pose, the gripper approaches the final grasp pose linearly. This approach allows for potential contact with the cloth but strictly avoids collisions with the robot itself or the environment. Upon reaching the actual grasp pose, the gripper is closed. Various distances were iteratively tested to determine a reachable and collision-free pregrasp pose.
@@ 3.4.1 Measuring Performance
The execution time of cloth manipulation algorithms is also an important metric to monitor Avigal et al. (2022); Garcia-Camacho et al. (2020) as some methods require far more interactions or employ elaborate recovery schemes. For our specific benchmark, which restricts the interactions with the cloth to a single grasp and stretch, we recommend setting a reasonable maximum execution time ( 30 seconds seems reasonable) for the grasp selection algorithm. In addition, the average execution can be reported, but this is optional.
@@ 3.4.2 Garments
We do not include a specific set of garments to test on for this benchmark, to avoid the additional effort of distributing such a physical dataset. We recommend ensuring diversity in terms of appearance, cloth materials and garment types as cloth manipulation algorithms should be able to generalize to these properties. If possible, we recommend to include towels from the Household Cloth Dataset Garcia-Camacho et al. (2022), which is a valuable effort to standardize cloth manipulation but lacks the visual diversity to properly test generalization. The garment set used for our competition can serve as inspiration.
For evaluating cloth manipulation, the initial configuration of the garment usually influences the algorithmic results significantly Gusseme et al. (2025), warranting a protocol Proesmans et al. (2023) to ensure the distribution of initial configurations is both appropriate and reproducible (as far is this is possible for clothes). However, this makes real-world evaluation even slower. In this benchmark, the initial configuration is less important as we use a two-phase initialization by grasping the highest and lowest point, as explained in Section 3.1. As such, to reduce the evaluation effort, we recommend to simply place the cloth manually on the surface in a random configuration before initializing the evaluation procedure. Nonetheless, the initial configuration will remain a source of randomness. We believe this is unavoidable for cloth benchmarking, as it is infeasible to reproduce a garment configuration entirely. To limit the impact, evaluators should ensure they make the initial configurations as random as possible and as similar as possible between different evaluations.
@@ 4.2 Data Collection
The annotators were PhD researchers from IDLab-AIRO, some specializing in robotic cloth manipulation, while others had expertise in unrelated areas. This diverse expertise resulted in varying degrees of success, exposing participants to both effective and ineffective grasps. We found it useful to provide the annotators with a virtual top-down view, obtained by reprojecting the point cloud, in addition to the actual camera viewpoint. They annotated the grasp position on the point cloud and the orientation was chosen to be tangential to the point cloud. Although the annotators used various strategies, many used their internal understanding of clothes and tried to locate semantic locations on the cloth that were visible and appropriate for the other grasp point. This is similar to the methods used by earlier work Hata et al. (2008); Maitin-Shepard et al. (2010); Doumanoglou et al. (2016). Different annotators had different risk-reward trade-offs when choosing to grasp partially occluded semantic locations rather than locations that would result in lower coverage but were easier to grasp.
@@
A limitation of the training set of the dataset is that a number of videos are incomplete due to a recording issue. However, all 176 live evaluation trials were fully recorded, ensuring the integrity of the competition data. These videos are provided for completeness but were not used for training by the competition participants. We also stress that the demonstrations are not perfect, they include trials with grasp failures and/or low coverage. We did not remove these as they could be useful in some contexts, but we encourage users of our dataset to filter the dataset to suit their needs.
@@ 5.2 Participant’s Grasp Selection Algorithms
The second learning-based method that focuses on graspability, Grasp-Cloth by Samsung Research China - Beijing starts by dividing the point cloud into key areas, for example around the bottom left corner and the bottom right corner for towels. These key areas were validated as good grasp regions by using the PyFleX cloth simulator. Subsequently, grasps for each region are generated using a grasp generation network that they trained on Graspnet-1Billion (Fang et al., 2020b), even though this dataset notably lacks cloth-like objects. This network is thus used purely to assess graspability in a geometric sense and has no notion of cloth behavior or the task of unfolding. This is also the only method that did not use a fixed grasp depth but instead used the grasp depth outputted from the neural network.
@@
Rank Team Method
Grasp
success rate
Coverage for
successful grasps
Coverage
@@
The CFAN (Cloth Flattening with Affordance learning and surface Normal estimation) method by AI&ROBOT LAB employs a neural network to generate an affordance map from depth images (Sunil et al., 2023). The point with the highest affordance is selected as the grasp location, and the surface normal at that point determines the grasp orientation. The Grasp with PointNet-VAE method by 3C1S utilizes a PointNet-like architecture that takes the segmented colored point clouds and outputs two points. The first point is used as grasp location and the second point which has to be adjacent to the first is used to determine grasp orientation. This pose prediction is then refined further using a variational autoencoder (VAE). The team manually filtered the provided dataset to remove data points with failed grasps or suboptimal coverage.
@@ 5.3 Competition Results
An even more detailed view of all 176 evaluation trial run at ICRA is given in Figure 7. Our analysis of all trials across all teams reveals several notable observations. Shirts had a significantly higher grasp success rate (0.7) than towels (0.53), likely attributable to their lower bending stiffness, which facilitates grasping. Despite this difference in grasp success, the average coverage achieved by successful grasps was similar for both shirts (0.60) and towels (0.57). Furthermore, no significant difference in average coverage was observed between seen (0.49) and unseen items (0.47), suggesting that most methods generalize well to unseen garments.
@@ 6 Discussion
When comparing the competition results to prior work, a noticeable gap emerges between the average coverage achieved here (0.6 by the top team) and values reported in previous studies (e.g., 0.8 by FlingBot (Ha and Song, 2022) and 0.85 by UnFoldIR (Proesmans et al., 2023)). However, this discrepancy can be attributed to several factors that make direct comparisons challenging. Subtle differences in experimental setups, such as the cloth’s initial configuration or the choice of evaluation items, can significantly influence unfolding difficulty. Moreover, prior work often reports coverage after multiple actions, recovery attempts, or filtering out grasp failures—all of which can substantially inflate performance metrics. Our competition’s live, single-attempt format, diversity in garments and automated evaluation minimize potential biases, offering a more realistic assessment of cloth unfolding capabilities under real-world constraints. This highlights the critical importance of independent evaluations outside controlled laboratory settings.
@@
The competitive performance of the second-placed method showcases the potential of learning-based approaches. It also illustrates the usefulness of our dataset for learning in-air unfolding and shows that such methods are able to generalize to different scenes and even unseen clothes, even when they use RGB images instead of geometric depth images as input. However, reaching their full potential likely requires even more high-quality data, which is expensive to collect on real robots. The competitiveness of traditional methods suggests they could play a key role in autonomous data collection, effectively serving to “jump-start the data flywheel”.
@@ 7 Conclusion
Julius Hietala is with Nimble Robotics Inc., San Francisco, USA.
@@
Ning Guo, Fang Wan, and Chaoyang Song are with the Southern University of Science and Technology, Shenzhen, China.

### v2

@@ Abstract
Robotic cloth manipulation suffers from a lack of standardized benchmarks and shared datasets for evaluating and comparing different approaches. To address this, we created a benchmark and organized the ICRA 2024 Cloth Competition, a unique head-to-head evaluation focused on grasp pose selection for in-air robotic cloth unfolding. Eleven teams participated in the competition, utilizing the publicly released dataset of 500 real-world robotic grasp attempts for cloth unfolding and employing diverse approaches to generate in-air unfolding grasps. Analysis of the competition results revealed insights about the trade-off between grasp success and coverage, the surprisingly strong achievements of hand-engineered methods and a significant discrepancy between competition performance and prior work, underscoring the importance of independent, out-of-the-lab evaluation in robotic cloth manipulation. We also expanded the dataset with 176 competition evaluation trials, resulting in a dataset of 679 unfolding demonstrations across 34 garments. This dataset is a valuable resource for developing and evaluating grasp selection methods, particularly for learning-based approaches. We hope that the benchmark, dataset and competition results can serve as a foundation for future benchmarks and drive further progress in data-driven robotic cloth manipulation.
@@ 2.1.1 Overview of Cloth Unfolding Strategies
Cloth unfolding is a critical prerequisite for various downstream cloth manipulation tasks such as folding, hanging, ironing, and assisted dressing. Researchers have explored various unfolding strategies, with computer vision often playing an important role. For a general introduction to using computer vision for cloth manipulation, we refer the reader to the surveys by Jiménez (2017) and Jiménez and Torras (2020). The strategy evaluated in the benchmark and competition aligns with what Jimenez identifies as Grasp Point Localisation, one of three key visual tasks he proposes as fundamental to cloth manipulation, alongside classification and state estimation. Grasp point localisation is deemed fundamental because, as Jiménez (2017) aptly states: “Any manipulative action requires instructing the robot where to grasp.” Combining well-placed grasps with open-loop motions can be sufficient for completing many manipulation tasks. For instance, unfolding might be achieved with a predetermined stretch motion after determining two appropriate grasp poses on the cloth.
@@
  • In-air grasping: Grasping a new point in the air while holding the cloth with one gripper and then stretching the cloth to unfold (Maitin-Shepard et al., 2010; Doumanoglou et al., 2014a). Starting from a crumpled garment, a common heuristic to bootstrap this strategy is first grasping its highest (or a random) point, lifting it, and then grasping its lowest point. This is the strategy used in the benchmark, and we also use the lowest point heuristic as the initialisation procedure.
  • Lifting and laying down: Simultaneous grasping of two points on a cloth lying on a surface, followed by lifting and laying it back down (Martínez and Ruiz-del-Solar, 2017; Thananjeyan et al., 2022). FlingBot exemplifies this approach, incorporating a flinging motion before laying down the cloth (Ha and Song, 2022).
@@
While not mutually exclusive (humans often seamlessly switch between strategies), these categories offer a framework for understanding the diverse approaches in the literature. A key commonality among the first three categories is the focus on achieving two “good” grasps on the cloth as a primary means of unfolding. These strategies operate under the generally valid assumption that grasping a cloth item at two suitable points can largely accomplish unfolding. Furthermore, due to the nature of the unfolding task, nearly all the aforementioned strategies can be applied iteratively, with the option to reset by simply releasing the cloth and starting over. This iterative nature necessitates that evaluation metrics like success rate or final coverage are always contextualized in the total amount of time taken, as even very inefficient strategies can eventually succeed through sufficient attempts.
@@ 2.2 Datasets for Robotic Cloth Manipulation
Datasets like Dex-Net 2.0 (Mahler et al., 2017) and GraspNet-1Billion (Fang et al., 2020b) were instrumental in advancing learning-based grasping, primarily for pick-and-place tasks with rigid objects. These works relied on analytical measures to determine grasp stability, focusing on feasibility and force closure. However, grasp selection for cloth unfolding presents a more complex challenge. The deformable nature of cloth makes it difficult to analytically determine grasp success, as the interaction between the gripper and the fabric can lead to unpredictable deformations. Additionally, the relative grasp location on the cloth significantly influences the unfolding outcome, making it crucial to consider not only grasp stability but also the subsequent manipulation trajectory. This distinction highlights the need for specialized cloth manipulation datasets.
@@ 3.2 Grasp Execution
To maximize both grasp success and workspace coverage while avoiding premature collisions, a two-part motion strategy is used for executing the provided grasps. This strategy consists of first moving the arm to a collision-free pregrasp pose, positioned a few centimeters behind the final grasp pose (i.e., the orientation is the same, but there is a small translation along the negative z-axis of the gripper, compared to the actual grasp frame). Then, from this pregrasp pose, the gripper approaches the final grasp pose linearly. This approach allows for potential contact with the cloth but strictly avoids collisions with the robot itself or the environment. Upon reaching the actual grasp pose, the gripper is closed. Various distances were iteratively tested to determine a reachable and collision-free pregrasp pose.
@@ 3.4.1 Measuring Performance
The execution time of cloth manipulation algorithms is also an important metric to monitor Avigal et al. (2022); Garcia-Camacho et al. (2020) as some methods require far more interactions or employ elaborate recovery schemes. For this benchmark, which restricts the interactions with the cloth to a single grasp and stretch, we recommend setting a reasonable maximum execution time (30 seconds seems reasonable) for the grasp selection algorithm. In addition, the average execution can be reported, but this is optional.
@@ 3.4.2 Garments
We do not include a specific set of garments to test on for this benchmark, to avoid the additional effort of distributing such a physical dataset. We recommend ensuring diversity in terms of appearance, cloth materials and garment types, as cloth manipulation algorithms should be able to generalize to these properties. If possible, we recommend including towels from the Household Cloth Dataset Garcia-Camacho et al. (2022), which is a valuable effort to standardize cloth manipulation but lacks the visual diversity to properly test generalization. The garment set used for this competition can serve as inspiration.
For evaluating cloth manipulation, the initial configuration of the garment usually influences the algorithmic results significantly Gusseme et al. (2025), warranting a protocol Proesmans et al. (2023) to ensure the distribution of initial configurations is both appropriate and reproducible (as far as this is possible for clothes). However, this makes real-world evaluation even slower. In this benchmark, the initial configuration is less important as we use a two-phase initialization by grasping the highest and lowest point, as explained in Section 3.1. As such, to reduce the evaluation effort, we recommend simply placing the cloth manually on the surface in a random configuration before initializing the evaluation procedure. Nonetheless, the initial configuration will remain a source of randomness. We believe this is unavoidable for cloth benchmarking, as it is infeasible to reproduce a garment configuration entirely. To limit the impact, evaluators should ensure they make the initial configurations as random as possible and as similar as possible between different evaluations.
@@ 4.2 Data Collection
The annotators were PhD researchers from IDLab-AIRO, some specializing in robotic cloth manipulation, while others had expertise in unrelated areas. This diverse expertise resulted in varying degrees of success, exposing participants to both effective and ineffective grasps. We found it useful to provide the annotators with a virtual top-down view, obtained by reprojecting the point cloud, in addition to the actual camera viewpoint. They annotated the grasp position on the point cloud, and the orientation was chosen to be tangential to the point cloud. Although the annotators used various strategies, many used their internal understanding of clothes and tried to locate semantic locations on the cloth that were visible and appropriate for the other grasp point. This is similar to the methods used by earlier work Hata et al. (2008); Maitin-Shepard et al. (2010); Doumanoglou et al. (2016). Different annotators had different risk-reward trade-offs when choosing to grasp partially occluded semantic locations rather than locations that would result in lower coverage but were easier to grasp.
@@
Note that some demonstrations in the dataset contain suboptimal grasp actions. The dataset (both train demonstrations and competition trials) includes trials with grasp failures (71/505) and/or low coverage. We did not remove these as they could be useful in some contexts, but we encourage users to filter the dataset to suit their needs. To facilitate filtering, the grasp success rate is available for each sample. Additionally, we provided a segmentation mask of the cloth in the final observation and a reference mask of the cloth item, which can be used to calculate coverage for each demonstration.
@@ 5.2 Participant’s Grasp Selection Algorithms
The second learning-based method that focuses on graspability, Grasp-Cloth by Samsung Research China - Beijing, starts by dividing the point cloud into key areas, for example, around the bottom left corner and the bottom right corner for towels. These key areas were validated as good grasp regions by using the PyFleX cloth simulator. Subsequently, grasps for each region are generated using a grasp generation network that they trained on Graspnet-1Billion (Fang et al., 2020a), even though this dataset notably lacks cloth-like objects. This network is thus used purely to assess graspability in a geometric sense and has no notion of cloth behavior or the task of unfolding. This is also the only method that did not use a fixed grasp depth but instead used the grasp depth outputted from the neural network.
@@
Rank Team Method Grasp success rate Coverage for successful grasps Coverage
@@
The CFAN (Cloth Flattening with Affordance learning and surface Normal estimation) method by AI&ROBOT LAB employs a neural network to generate an affordance map from depth images (Sunil et al., 2023). The point with the highest affordance is selected as the grasp location, and the surface normal at that point determines the grasp orientation. The Grasp with PointNet-VAE method by 3C1S utilizes a PointNet-like architecture that takes the segmented colored point clouds and outputs two points. The first point is used as grasp location and the second point, which has to be adjacent to the first, is used to determine grasp orientation. This pose prediction is then refined further using a variational autoencoder (VAE). The team manually filtered the provided dataset to remove data points with failed grasps or suboptimal coverage.
@@ 5.3 Competition Results
An even more detailed view of all 176 evaluation trial run at ICRA is given in Figure 7. The analysis of all trials across all teams reveals several notable observations. Shirts had a significantly higher grasp success rate (0.7) than towels (0.53), likely attributable to their lower bending stiffness, which facilitates grasping. Despite this difference in grasp success, the average coverage achieved by successful grasps was similar for both shirts (0.60) and towels (0.57). Furthermore, no significant difference in average coverage was observed between seen (0.49) and unseen items (0.47), suggesting that most methods generalize well to unseen garments.
@@ 6 Discussion
When comparing the competition results to prior work, a noticeable gap emerges between the average coverage achieved here (0.6 by the top team) and values reported in previous studies (e.g., 0.8 by FlingBot (Ha and Song, 2022) and 0.85 by UnFoldIR (Proesmans et al., 2023)). However, this discrepancy can be attributed to several factors that make direct comparisons challenging. Subtle differences in experimental setups, such as the cloth’s initial configuration or the choice of evaluation items, can significantly influence unfolding difficulty. Moreover, prior work often reports coverage after multiple actions, recovery attempts, or filtering out grasp failures—all of which can substantially inflate performance metrics. The competition’s live, single-attempt format, diversity in garments and automated evaluation minimize potential biases, offering a more realistic assessment of cloth unfolding capabilities under real-world constraints. This highlights the critical importance of independent evaluations outside controlled laboratory settings.
@@
The competitive performance of the second-placed method showcases the potential of learning-based approaches. It also illustrates the usefulness of the dataset for learning in-air unfolding and shows that such methods are able to generalize to different scenes and even unseen clothes, even when they use RGB images instead of geometric depth images as input. However, reaching their full potential likely requires even more high-quality data, which is expensive to collect on real robots. The competitiveness of traditional methods suggests they could play a key role in autonomous data collection, effectively serving to “jump-start the data flywheel”.
@@ 7 Conclusion
Julius Hietala is an independent researcher
@@
Ning Guo and Fang Wan are with the Southern University of Science and Technology, Shenzhen, China, supported by the National Natural Science Foundation of China under Grant 62206119 and Shenzhen Long-Term Support for Higher Education at SUSTech under Grant 20231115141649002. Chaoyang Song is with the Mohamed bin Zayed University of Artificial Intelligence, Abu Dhabi, United Arab Emirates, supported by the National Natural Science Foundation of China under Grant 62473189.